SpaceRef

SpaceRef


Aqua Weekly Report 31 May 2002 - 7 Jun 2002

Status Report From: Goddard Space Flight Center
Posted: Friday, June 7, 2002

Checkout of the Aqua spacecraft continues to go extremely well. We completed our second (of six) Ascent Maneuver nominally on Wednesday (6/5) and plan to perform the third tomorrow (6/8) morning.

Last Friday (5/31), we performed a series of Thruster Characterization tests. The performance of thrusters 3 & 4 agreed well with their predicted operation from simulations. Thrusters 1 & 2 showed discrepancies in pitch & roll from their predicts which were attributed to plume effects. The GNC subsystem team modeled these effects into their simulation and generated a FSW patch with new steady state K-parameters. This patch was uploaded prior to Ascent Maneuver #2 and the new K-parameters improved the burn duty cycle and attitude/rate transients. GNC believes that the Center-of-Gravity (CG), moment of inertia, and thruster alignments are all nominal.

Instrument checkout activities also continued nominally throughout the week. The instrument status is:

  • AIRS: Currently in Decontaminate mode. Other than parking the scanner for the Ascent Maneuver, AIRS had no planned activities this past week. They plan to turn off their decontaminate heater and initialize/begin cooler operations next week.
  • AMSR-E: Currently in Normal mode, rotating at 40 rpm. AMSR-E has completed their activation and checkout activities. Last week, AMSR-E adjusted their Automatic Gain Control (AGC) level resulting in improved data images. No further instances of the SPS anomaly occurred last week and the investigation into the cause of the anomaly continues.
  • AMSU-A: Currently in Full Scan mode. Other than parking the instruments for the Ascent Maneuver, AMSU has no further scheduled activities until we complete the ascent burns to our 705km mission altitude, currently planned for mid/late June. Several IOT training exercises were performed last week.
  • CERES: Currently in Aft- Biaxial, Fore - Crosstrack mode with their covers closed. CERES nominally completed several checkout activities this past week, including the Fore instrument Biaxial Scan Rate Tests, enabling of the Fore/Aft Solar Presence Sensors (SPSs), and enabling of the CERES MCL expire TMON. They also stopped scanning for the Ascent Maneuver. Several CERES operations are now being controlled from their MCL, including their solar avoidance activities. They plan to open their covers after we complete the Ascent Maneuvers.
  • HSB: Currently in Scan Normal mode. Other than parking the instrument for the Ascent Maneuver, HSB, like AMSU-A, has no further scheduled activities until we attain our mission altitude. Several IOT training exercises were performed last week.
  • MODIS: Currently in Standby mode, with the Nadir and Solar Diffuser doors unlatched and closed, and the Space View Door (SVD) unlatched and at the OPEN position. MODIS uplinked their SCS loads early last week and turned off their outgas heater (today). They plan transition to science mode later today (6/7) and perform their first Photovoltaic Electronic Calibration (PV ECal).
Today (6/7), we also began characterization of our TDRSS capability at 8k and 16k telemetry rates. Nominally, we utilize a 4k maximum telemetry rate through TDRSS, however, data at the 4k rate indicated that we may have the signal margin at the greater rates. Both tests proved successful. We were able to lock-up at both 8k and 16k. Link margins remained high and bit error rates remained low. Future field-of-view tests will be scheduled.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.